# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 2.1.2 Partial Contraction
Consider the following nonlinear system with the state $x\in\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ and the auxiliary or virtual system with the state $q\in\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$:
@@
The theorem statement follows from Theorem 2.1 and the fact that $q=x$ and a trajectory with the specific property are particular solutions of (2.17) (see Wang and Slotine [2005] for details). $\qed$
The importance of this theorem lies in the fact that we can analyze contraction of some specific parts of the system (2.16) while treating the rest as a function of the time-varying parameter $x(t)$. Strictly speaking, the system (2.16) is said to be partially contracting, but we will not distinguish partial contraction from contraction of Definition 2.3 in this paper for simplicity. Instead, we will use the variable $q$ when referring to partial contraction of Theorem 2.2.
@@
Let us illustrate the role of partial contraction using the following nonlinear system Jouffroy and Slotine [2004]:
@@
which also has $q=x$ and $q=x_{d}$ as its particular solutions, proving contraction is no longer straightforward because of the terms $\partial D/\partial q_{i}$ and $\partial K/\partial q_{i}$ in the differential dynamics of (2.21). This is due to the fact that, in contrast to (2.20), (2.21) has particular solutions nonlinear in $q$ in addition to $q=x$ and $q=x_{d}$, and the condition (2.6) becomes more involved for (2.21) as it is for guaranteeing exponential convergence of any couple of these particular solution trajectories Jouffroy and Slotine [2004].
@@
As one of the key applications of partial contraction given in Theorem 2.2, let us consider the following closed-loop Lagrangian system [Slotine and Li, 1991, p. 392]:
@@
where the skew-symmetry of $\dot{\mathcal{H}}-2\mathcal{C}$ is used to obtain the second equality. Since $\mathcal{K}\succ 0$, (2.25) indicates that the virtual system (2.24) is partially contracting in $q$ with $\mathcal{H}$ defining its contraction metric. Contraction of the full state $(\mathtt{q},\dot{\mathtt{q}})$ will be discussed in Example 2.7.
Note that if we treat the arguments $(\mathtt{q},\dot{\mathtt{q}})$ of $\mathcal{H}$ and $\mathcal{C}$ also as the virtual state $q$, we end up having additional terms such as $\partial\mathcal{H}/\partial{q_{i}}$, which makes proving contraction analytically more involved as in Example 2.5.
As can be seen from Examples 2.5 and 2.6, the role of partial contraction in theorem 2.2 is to provide some insight on stability even for cases where it is difficult to prove contraction for all solution trajectories as in Theorem 2.1. Although finding a contraction metric analytically for general nonlinear systems is challenging, we will see in Sec. 3 and Sec. 4 that the convex nature of the contraction condition (2.6) helps us find it numerically.
2.2 Path-Length Integral and Robust Incremental Stability Analysis
Theorem 2.1 can also be proven by using the transformed squared length integrated over two arbitrary solutions of (2.1Lohmiller and Slotine [1998], Chung et al. [2013], Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019], which enables formalizing its connection to incremental stability discussed in Definition 2.3. Note that the integral forms (2.26) and (2.27) to be given in Theorem 2.3 are useful for handling perturbed systems with external disturbances as shall be seen in Theorems 2.42.8.
@@ 2.2 Path-Length Integral and Robust Incremental Stability Analysis
Let $\xi_{0}$ and $\xi_{1}$ be the two arbitrary solutions of (2.1), and define the transformed squared length with $M(x,t)$ of Theorem 2.1 as follows:
@@
where $M(x,t)\succeq\underline{m}\mathrm{I},~\forall x,t$ for $\exists\underline{m}\in\mathbb{R}_{>0}$, and Theorem 2.1 can also be proven by using (2.26) and (2.27) as a Lyapunov-like function, resulting in incremental exponential stability of the system (2.1) (see Definition 2.2). Note that the shortest path integral $V_{\ell}$ of (2.27) with a parameterized state $x$ (i.e., $\inf V_{\ell}=\sqrt{\inf V_{s\ell}}$) defines the Riemannian distance and the path integral of a minimizing geodesic Manchester and Slotine [2017].
@@
Using $M(x,t)\succeq\underline{m}\mathrm{I}$ which gives $\sqrt{\underline{m}}\|\xi_{1}-\xi_{0}\|\leq V_{\ell}$, we have $\|\xi_{1}-\xi_{0}\|=\|\int^{\xi_{1}}_{\xi_{0}}\delta{x}\|\leq{V_{\ell}}/{\sqrt{\underline{m}}}$ of (2.28). The inequality $V_{\ell}\leq\sqrt{V_{s\ell}}$ of (2.28) can be proven by applying the Cauchy–Schwarz inequality [Horn and Johnson, 2012, p. 316] to the functions $\psi_{1}(\mu)=\|\Theta(x,t)(\partial x/\partial\mu)\|$ and $\psi_{2}(\mu)=1$.
@@
to have $\dot{V}_{s\ell}\leq-2\alpha V_{s\ell}$ and $\dot{V}_{\ell}\leq-\alpha V_{\ell}$ by the contraction conditions (2.5) and (2.6). Since these hold for any $\xi_{0}$ and $\xi_{1}$, the incremental exponential stability results in Theorem 2.1 follow from the comparison lemma of Lemma 2.1 (see also Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019], and Manchester and Slotine [2017], Singh et al. [2017] for the discussion on the geodesic). $\qed$
2.2.1 Deterministic Perturbation
@@
where $d_{\mu}(\mu,\xi_{1},t)=\mu d(\xi_{1},t)$. Since contraction means exponential convergence, a contracting system exhibits a superior property of robustness Lohmiller and Slotine [1998], Chung et al. [2013].
@@
If the system (2.1) satisfies (2.5) and (2.6) of Theorem 2.1 (i.e., the system (2.1) is contracting), then the path integral $V_{\ell}(q,\delta q,t)=\int^{\xi_{1}}_{\xi_{0}}\|\Theta(q,t)\delta q\|$ of (2.27), where $\xi_{0}$ is a solution of the contracting system (2.1), $\xi_{1}$ is a solution of the perturbed system (2.30), and $q$ is the virtual state of (2.31), exponentially converges to a bounded error ball as long as ${\Theta}{d}\in\mathcal{L}_{\infty}$ (i.e., $\sup_{x,t}\|\Theta d\|<\infty$). Specifically, if $\exists\underline{m},\overline{m}\in\mathbb{R}_{>0}$ and $\exists\bar{d}\in\mathbb{R}_{\geq 0}$ s.t. $\bar{d}=\sup_{{x},t}\|{d}(x,t)\|$ and
@@
Using the contraction condition (2.6), we have for $M=\Theta^{\top}\Theta$ given in Theorem 2.1 that
@@
which implies $\dot{V}_{\ell}(t)\leq-\alpha V_{\ell}(t)+\sup_{q,\xi_{1},t}{\|{\Theta}({q},t)d(\xi_{1},t)\|}$ for $V_{\ell}$ in (2.27) of Theorem 2.3. Thus, applying the comparison lemma (see Lemma 2.1) results in
@@ 2.2.2 Stochastic Perturbation
where $G:\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times w}$ is a matrix-valued function and $\mathscr{W}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{w}$ is a $w$-dimensional Wiener process [Arnold, 1974, p. 100] (see also [Arnold, 1974, p. xii] for the notations used). For the sake of the existence and uniqueness of the solution, we assume in (2.37) that
@@
In order to analyze the incremental stability property of (2.37) as in Theorem 2.3, we consider two trajectories $\xi_{0}(t)$ and $\xi_{1}(t)$ of stochastic nonlinear systems with Gaussian white noise, driven by two independent Wiener processes $\mathscr{W}_{0}(t)$ and $\mathscr{W}_{1}(t)$:
@@
One can show that (2.37) has a unique solution $x(t)$ which is continuous with probability one under the conditions (2.38) and (2.39) (see [Arnold, 1974, p. 105] and Pham et al. [2009], Dani et al. [2015]), leading to the following lemma as in the comparison lemma of Lemma 2.1.
@@
where $\gamma\in\mathbb{R}_{>0}$, $c\in\mathbb{R}_{\geq 0}$, and $\mathscr{L}$ denotes the infinitesimal differential generator of the Itô process given in [Kushner, 1967, p. 15]. Then we have the following bound Pham et al. [2009]:
@@
The bound (2.42) follows from Theorem 2 of Pham et al. [2009] (see also Tsukamoto and Chung [2021], Dani et al. [2015], Tarn and Rasis [1976], Zakai [1967], Tsukamoto and Chung [2019] and [Kushner, 1967, p. 10] (Dynkin’s formula)). The probability tracking error bound (2.43) then follows from Markov’s inequality [Geoffrey R. Grimmett, 2001, pp. 311-312]. $\qed$
@@
Although Lemma 2.2 considers the second moment of $\|\xi_{1}(t)-\xi_{0}(t)\|$, i.e., $\mathbb{E}[\|\xi_{1}(t)-\xi_{0}(t)\|^{2}]$, it can be readily generalized to the $p$-th moment of $\|\xi_{1}(t)-\xi_{0}(t)\|$, i.e., $\mathbb{E}[\|\xi_{1}(t)-\xi_{0}(t)\|^{p}]$, applying the Lyapunov-based technique proposed in Mazumdar et al. [2020].
@@
where $G(\mu,\xi_{0},\xi_{1},t)=[(1-\mu)G_{0}(\xi_{0},t),\mu G_{1}(\xi_{1},t)]$ and $\mathscr{W}=[\mathscr{W}_{0}^{\top},\mathscr{W}_{1}^{\top}]^{\top}$. As a consequence of Lemma 2.2, showing stochastic incremental stability between $\xi_{0}$ and $\xi_{1}$ of (2.40) reduces to proving the relation (2.41), similar to the deterministic case in Theorems 2.1 and 2.4.
@@
Suppose that $\exists\bar{g}_{0}\in\mathbb{R}_{\geq 0}$ and $\exists\bar{g}_{1}\in\mathbb{R}_{\geq 0}$ s.t. $\sup_{x,t}\|G_{1}(x,t)\|_{F}=\bar{g}_{0}$ and $\sup_{x,t}\|G_{1}(x,t)\|_{F}=\bar{g}_{1}$ in (2.40). Suppose also that there exists $M(x,t)\succ 0,~\forall x,t$, s.t. $M_{x_{i}}=\partial M/\partial x_{i}$ is Lipschitz with respect to $x$ for all $i=1,\cdots,n$, i.e., $\exists L_{m}\in\mathbb{R}_{\geq 0}$ s.t.
@@
By definition of the infinitesimal differential generator given in Lemma 2.2 [Kushner, 1967, p. 15], we have Tsukamoto and Chung [2021, 2019]
@@
Since $M_{x_{i}}$ is Lipschitz as in (2.45), we have $\|M_{x_{i}x_{j}}\|\leq L_{m}$ and $\left\|M_{x_{i}}\right\|\leq\sqrt{{2L_{m}}{\overline{m}}}$ using (2.32) as derived Tsukamoto et al. [2021]. Computing $\mathscr{L}V_{s\ell}$ of (2.49) using these bounds, the bounds of $\|G_{0}\|_{F}$ and $\|G_{1}\|_{F}$, and $\partial_{\mu}G=\partial G/\partial\mu=[-G_{0},G_{1}]$ as in Tsukamoto and Chung [2021, 2019] yields
@@
Although we consider the Gaussian white noise stochastic differential equation (2.40) when referring to stochastic systems in this paper, other types of stochastic noises, including compound Poisson shot noise and bounded-measure Lévy noise, could be considered as in Theorem 2.5 using contraction theory Han and Chung [2021].
2.3 Finite-Gain $\mathcal{L}_{p}$ Stability and Contraction of Hierarchically Combined Systems
Due to the simpler stability analysis of Theorem 2.1 when compared with Lyapunov theory, Input-to-State Stability (ISS) and input-output stability in the sense of finite-gain $\mathcal{L}_{p}$ stability can be easily studied using contraction theory Chung et al. [2013].

### v8

@@ 2.1.2. Partial Contraction
Partial ContractionThm:partial_contraction Consider the following nonlinear system with the state $x\in\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ and the auxiliary or virtual system with the state $q\in\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$:
@@
This statement follows from Theorem LABEL:THM:Thm:contraction and the fact that $q=x$ and a trajectory with the specific property are particular solutions of (2.17) (see Ref:contraction3 for details). $\qed$
The importance of this theorem lies in the fact that we can analyze contraction of some specific parts of the system (2.16) while treating the rest as a function of the time-varying parameter $x(t)$. Strictly speaking, the system (2.16) is said to be partially contracting, but we will not distinguish partial contraction from contraction of Definition LABEL:DEF:Def:contraction in this paper for simplicity. Instead, we will use the variable $q$ when referring to partial contraction of Theorem LABEL:THM:Thm:partial_contraction.
@@
Partial Contraction Iex:partial_cont Let us illustrate the role of partial contraction using the following nonlinear system Ref:contraction6:
@@
which also has $q=x$ and $q=x_{d}$ as its particular solutions, proving contraction is no longer straightforward because of the terms $\partial D/\partial q_{i}$ and $\partial K/\partial q_{i}$ in the differential dynamics of (2.21). This is due to the fact that, in contrast to (2.20), (2.21) has particular solutions nonlinear in $q$ in addition to $q=x$ and $q=x_{d}$, and the condition (2.6) becomes more involved for (2.21) as it is for guaranteeing exponential convergence of any couple of these particular solution trajectories Ref:contraction6.
@@
Partial Contraction IIex:lag_metric As one of the key applications of partial contraction given in Theorem LABEL:THM:Thm:partial_contraction, let us consider the following closed-loop Lagrangian system [Ref_Slotine, p. 392]:
@@
where the skew-symmetry of $\dot{\mathcal{H}}-2\mathcal{C}$ is used to obtain the second equality. Since $\mathcal{K}\succ 0$, (2.25) indicates that the virtual system (2.24) is partially contracting in $q$ with $\mathcal{H}$ defining its contraction metric. Contraction of the full state $(\mathtt{q},\dot{\mathtt{q}})$ will be discussed in Example LABEL:EX:ex:hiera.
Note that if we treat the arguments $(\mathtt{q},\dot{\mathtt{q}})$ of $\mathcal{H}$ and $\mathcal{C}$ also as the virtual state $q$, we end up having additional terms such as $\partial\mathcal{H}/\partial{q_{i}}$, which makes proving contraction analytically more involved as in Example LABEL:EX:ex:partial_cont.
As can be seen from Examples LABEL:EX:ex:partial_cont and LABEL:EX:ex:lag_metric, the role of partial contraction in theorem LABEL:THM:Thm:partial_contraction is to provide some insight on stability even for cases where it is difficult to prove contraction for all solution trajectories as in Theorem LABEL:THM:Thm:contraction. Although finding a contraction metric analytically for general nonlinear systems is challenging, we will see in Sec. 3 and Sec. 4 that the convex nature of the contraction condition (2.6) helps us find it numerically.
2.2. Path-Length Integral and Robust Incremental Stability Analysis
Theorem LABEL:THM:Thm:contraction can also be proven by using the transformed squared length integrated over two arbitrary solutions of (2.1Ref:contraction1, Ref:phasesync, mypaperTAC, Ref:Stochastic, mypaper, which enables formalizing its connection to incremental stability discussed in Definition LABEL:DEF:Def:contraction. Note that the integral forms (2.26) and (2.27) to be given in Theorem LABEL:THM:Thm:path_integral are useful for handling perturbed systems with external disturbances as shall be seen in Theorems LABEL:THM:Thm:Robust_contraction_originalLABEL:THM:Thm:discrete_contraction.
@@ 2.2. Path-Length Integral and Robust Incremental Stability Analysis
Path IntegralThm:path_integral Let $\xi_{0}$ and $\xi_{1}$ be the two arbitrary solutions of (2.1), and define the transformed squared length with $M(x,t)$ of Theorem LABEL:THM:Thm:contraction as follows:
@@
where $M(x,t)\succeq\underline{m}\mathrm{I},~\forall x,t$ for $\exists\underline{m}\in\mathbb{R}_{>0}$, and Theorem LABEL:THM:Thm:contraction can also be proven by using (2.26) and (2.27) as a Lyapunov-like function, resulting in incremental exponential stability of the system (2.1) (see Definition LABEL:DEF:Def:incremental). Note that the shortest path integral $V_{\ell}$ of (2.27) with a parameterized state $x$ (i.e., $\inf V_{\ell}=\sqrt{\inf V_{s\ell}}$) defines the Riemannian distance and the path integral of a minimizing geodesic ccm.
@@
Using $M(x,t)\succeq\underline{m}\mathrm{I}$ which gives $\sqrt{\underline{m}}\|\xi_{1}-\xi_{0}\|\leq V_{\ell}$, we have $\|\xi_{1}-\xi_{0}\|=\|\int^{\xi_{1}}_{\xi_{0}}\delta{x}\|\leq{V_{\ell}}/{\sqrt{\underline{m}}}$ of (2.28). The inequality $V_{\ell}\leq\sqrt{V_{s\ell}}$ of (2.28) can be proven by applying the Cauchy–Schwarz inequality [10.5555/2422911, p. 316] to the functions $\psi_{1}(\mu)=\|\Theta(x,t)(\partial x/\partial\mu)\|$ and $\psi_{2}(\mu)=1$.
@@
to have $\dot{V}_{s\ell}\leq-2\alpha V_{s\ell}$ and $\dot{V}_{\ell}\leq-\alpha V_{\ell}$ by the contraction conditions (2.5) and (2.6). Since these hold for any $\xi_{0}$ and $\xi_{1}$, the incremental exponential stability results in Theorem LABEL:THM:Thm:contraction follow from the comparison lemma of Lemma 2.1 (see also mypaperTAC, Ref:Stochastic, mypaper, and ccm, 7989693 for the discussion on the geodesic). $\qed$
2.2.1. Deterministic Perturbation
@@
where $d_{\mu}(\mu,\xi_{1},t)=\mu d(\xi_{1},t)$. Since contraction means exponential convergence, a contracting system exhibits a superior property of robustness Ref:contraction1, Ref:phasesync.
@@
Robust Deterministic ContractionThm:Robust_contraction_original If the system (2.1) satisfies (2.5) and (2.6) of Theorem LABEL:THM:Thm:contraction (i.e., the system (2.1) is contracting), then the path integral $V_{\ell}(q,\delta q,t)=\int^{\xi_{1}}_{\xi_{0}}\|\Theta(q,t)\delta q\|$ of (2.27), where $\xi_{0}$ is a solution of the contracting system (2.1), $\xi_{1}$ is a solution of the perturbed system (2.30), and $q$ is the virtual state of (2.31), exponentially converges to a bounded error ball as long as ${\Theta}{d}\in\mathcal{L}_{\infty}$ (i.e., $\sup_{x,t}\|\Theta d\|<\infty$). Specifically, if $\exists\underline{m},\overline{m}\in\mathbb{R}_{>0}$ and $\exists\bar{d}\in\mathbb{R}_{\geq 0}$ s.t. $\bar{d}=\sup_{{x},t}\|{d}(x,t)\|$ and
@@
Using the contraction condition (2.6), we have for $M=\Theta^{\top}\Theta$ given in Theorem LABEL:THM:Thm:contraction that
@@
which implies $\dot{V}_{\ell}(t)\leq-\alpha V_{\ell}(t)+\sup_{q,\xi_{1},t}{\|{\Theta}({q},t)d(\xi_{1},t)\|}$ for $V_{\ell}$ in (2.27) of Theorem LABEL:THM:Thm:path_integral. Thus, applying the comparison lemma (see Lemma 2.1) results in
@@ 2.2.2. Stochastic Perturbation
where $G:\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times w}$ is a matrix-valued function and $\mathscr{W}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{w}$ is a $w$-dimensional Wiener process [arnold_SDE, p. 100] (see also [arnold_SDE, p. xii] for the notations used). For the sake of the existence and uniqueness of the solution, we assume in (2.37) that
@@
In order to analyze the incremental stability property of (2.37) as in Theorem LABEL:THM:Thm:path_integral, we consider two trajectories $\xi_{0}(t)$ and $\xi_{1}(t)$ of stochastic nonlinear systems with Gaussian white noise, driven by two independent Wiener processes $\mathscr{W}_{0}(t)$ and $\mathscr{W}_{1}(t)$:
@@
One can show that (2.37) has a unique solution $x(t)$ which is continuous with probability one under the conditions (2.38) and (2.39) (see [arnold_SDE, p. 105] and Pham2009, Ref:Stochastic), leading to the following lemma as in the comparison lemma of Lemma 2.1.
@@
where $\gamma\in\mathbb{R}_{>0}$, $c\in\mathbb{R}_{\geq 0}$, and $\mathscr{L}$ denotes the infinitesimal differential generator of the Itô process given in [sto_stability_book, p. 15]. Then we have the following bound Pham2009:
@@
The bound (2.42) follows from Theorem 2 of Pham2009 (see also mypaperTAC, Ref:Stochastic, old_observer, doi:10.1137/0305037, mypaper and [sto_stability_book, p. 10] (Dynkin’s formula)). The probability tracking error bound (2.43) then follows from Markov’s inequality [probbook, pp. 311-312]. $\qed$
@@
Although Lemma 2.2 considers the second moment of $\|\xi_{1}(t)-\xi_{0}(t)\|$, i.e., $\mathbb{E}[\|\xi_{1}(t)-\xi_{0}(t)\|^{2}]$, it can be readily generalized to the $p$-th moment of $\|\xi_{1}(t)-\xi_{0}(t)\|$, i.e., $\mathbb{E}[\|\xi_{1}(t)-\xi_{0}(t)\|^{p}]$, applying the Lyapunov-based technique proposed in 9304491.
@@
where $G(\mu,\xi_{0},\xi_{1},t)=[(1-\mu)G_{0}(\xi_{0},t),\mu G_{1}(\xi_{1},t)]$ and $\mathscr{W}=[\mathscr{W}_{0}^{\top},\mathscr{W}_{1}^{\top}]^{\top}$. As a consequence of Lemma 2.2, showing stochastic incremental stability between $\xi_{0}$ and $\xi_{1}$ of (2.40) reduces to proving the relation (2.41), similar to the deterministic case in Theorems LABEL:THM:Thm:contraction and LABEL:THM:Thm:Robust_contraction_original.
@@
Robust Stochastic ContractionThm:robuststochastic Suppose that $\exists\bar{g}_{0}\in\mathbb{R}_{\geq 0}$ and $\exists\bar{g}_{1}\in\mathbb{R}_{\geq 0}$ s.t. $\sup_{x,t}\|G_{1}(x,t)\|_{F}=\bar{g}_{0}$ and $\sup_{x,t}\|G_{1}(x,t)\|_{F}=\bar{g}_{1}$ in (2.40). Suppose also that there exists $M(x,t)\succ 0,~\forall x,t$, s.t. $M_{x_{i}}=\partial M/\partial x_{i}$ is Lipschitz with respect to $x$ for all $i=1,\cdots,n$, i.e., $\exists L_{m}\in\mathbb{R}_{\geq 0}$ s.t.
@@
By definition of the infinitesimal differential generator given in Lemma 2.2 [sto_stability_book, p. 15], we have mypaperTAC, mypaper
@@
Since $M_{x_{i}}$ is Lipschitz as in (2.45), we have $\|M_{x_{i}x_{j}}\|\leq L_{m}$ and $\left\|M_{x_{i}}\right\|\leq\sqrt{{2L_{m}}{\overline{m}}}$ using (2.32) as derived nscm. Computing $\mathscr{L}V_{s\ell}$ of (2.49) using these bounds, the bounds of $\|G_{0}\|_{F}$ and $\|G_{1}\|_{F}$, and $\partial_{\mu}G=\partial G/\partial\mu=[-G_{0},G_{1}]$ as in mypaperTAC, mypaper yields
@@
Although we consider the Gaussian white noise stochastic differential equation (2.40) when referring to stochastic systems in this paper, other types of stochastic noises, including compound Poisson shot noise and bounded-measure Lévy noise, could be considered as in Theorem LABEL:THM:Thm:robuststochastic using contraction theory han2021incremental.
2.3. Finite-Gain $\mathcal{L}_{p}$ Stability and Hierarchical Contraction
Due to the simpler stability analysis of Theorem LABEL:THM:Thm:contraction when compared with Lyapunov theory, Input-to-State Stability (ISS) and input-output stability in the sense of finite-gain $\mathcal{L}_{p}$ stability can be easily studied using contraction theory Ref:phasesync.